# Explicit expressions for the gamma vector leading to connections to upper/lower bounds and structural properties

### v2

@@ Abstract
We find an explicit formula for the gamma vector in terms of the input polynomial in a way that extends it to arbitrary polynomials. More specifically, we find explicit linear combination in terms of coefficients of the input polynomial (using Catalan numbers and binomial coefficients) and an expression involving the derivative of the input polynomial. In the case where the input is the $h$-polynomial of a simplicial complex, this gives an interpretation of the gamma vector as a measure of differences in local and global contributions. We also apply them to connect signs/inequalities of (shifts of) the gamma vector to upper/lower bound conditions on coefficients of the input polynomial. Finally, we make use of the shape of the sums used to make these estimates and connections with intersection numbers to relate these properties of the gamma vector to algebraic structures.
@@ Introduction
Alternatively, it can be taken as the coefficients $\gamma_{i}$ used to write $h(t)$ as a linear combination of $t^{i}(1+t)^{n-2i}$ when $h(t)$ is a reciprocal polynomial. It was originally defined by Gal [16] in connection to real-rootedness questions of $h$-polynomials of flag spheres and had lower bound results with $h$-vectors in mind related to postivity of the gamma vector. Properties of this vector have since been involved in many different examples including those far beyond the original context [6].
Since the gamma vector of a polynomial is defined in terms of a functional equation involving the starting polynomial, it is natural to ask if there are explicit expressions in terms of coefficients of the original polynomial. We find explicit formulas for the gamma vector in terms of the coefficients of $h(t)$ in a way that complements the connections to lower bounds and derivatives involving the input polynomial mentioned above. In addition, they can be used to extend the definition of the gamma vector to arbitrary polynomials. Concretely, the components of the gamma vector/coefficients of $\gamma(u)$ written as a linear combination of the coefficients of $h(t)$ have coefficients coming from products of Catalan numbers and binomial coefficients (Theorem 1.1).
@@
Continuing in this direction, we also consider interactions with algebraic structures. Since many of the examples we had considered earlier were connected to independence complexes and broken circuits of matroids (which can be express in terms of intersection numbers), we were motivated to consider an input polynomial with coefficients coming from intersections of divsiors on a variety. So, we considered the case where the input polynomial is the specialization $V(u)=(Au+B)^{d}$ of the volume polynomial associated to a pair of divsiors $A,B$ on a smooth projective variety. In general, the gamma vector has the same sign as $(-1)^{r}$ when the intersections $A^{k}B^{d-k}$ alternate in sign. Suppose that these intersections are nonnegative. When we are working with $\mathbb{P}^{n}$, we can relate the sign of the gamma vector to the relative “size” of $A$ compared to $B$ (since these correspond to integers) (Example 3.4. When the ratio is very large, the gamma vector is negative. If it is very small, it is alternating in sign like the negative ratio case. Positivity properties of divisors are also related to the bounds in a more general setting since the $A^{k}B^{d-k}$ form a log concave sequence when $A$ and $B$ are nef. This can be used to reduce upper and lower bound conditions related to alternating sums of positive multiples of such terms to those on the first or last pair of elements in a sequence. Note that the connections to upper/lower bounds studied in the previous section carry over to volume polynomials with some additional interpretations and simplifications. Finally, we give an explanation of the $\mathbb{P}^{n}$ case in terms of Segre classes of vector bundles and make use of these classes to connect the rational functions giving components of the gamma vector in the derivative expression to positivity questions related to Segre classes (Example 3.7).
@@ Acknowledgments
I’d like to thank Lutz Warnke for pointing out a more recent reference [27] describing an elementary approach to the Lagrange inversion formula. This work was supported by Horizon Europe ERC Grant number: 101045750 / Project acronym: HodgeGeoComb.
@@ 1.1 An explicit formula involving Catalan numbers, binomial coefficients, and derivatives
[…]
@@ 2.1 Bounds involving translations
in order to have $\gamma_{r}\leq 0$ for even $r$ as well. This is a sort of lower bound condition. ∎
@@ 2.2 Upper/lower bounds and general gamma vectors
where we used the identity $\binom{p}{q}-\binom{p-1}{q-1}=\binom{p-1}{q}$ in the last line. This means that
@@
This follows directly from clearing denominators since $k(r-k-1)=kr-k^{2}-k\leq(k+1)(r-k)=kr-k^{2}+r-k$. The proof of Part 2a) implies that we have shown the statement of Part 2b) for $0\leq r\leq\frac{d}{3}$. To look at the remaining indices, we can split the first sum in Proposition 2.4 into those where $k\geq\frac{r}{d-2r}$ (decreasing coefficients of $b_{k}$) and $k<\frac{r}{d-2r}$ (increasing coefficients of $b_{k}$). The terms that we take alternating sums of are increasing on each part. Lemma 2.5 then shows that the terms have sign $(-1)^{r-1}$ and we have the claimed inequality. ∎

### v3

@@ Abstract
We find an explicit formula for the gamma vector in terms of the input polynomial in a way that extends it to arbitrary polynomials. More specifically, we find explicit linear combination in terms of coefficients of the input polynomial (using Catalan numbers and binomial coefficients) and an expression involving the derivative of the input polynomial. The first expression suggests connections to common Coxeter group/noncrossing partition structures in existing gamma positivity examples. In the case where the input is the $h$-polynomial of a simplicial complex, this gives an interpretation of the gamma vector as a measure of differences in local and global contributions. We also apply them to connect signs/inequalities of (shifts of) the gamma vector to upper/lower bound conditions on coefficients of the input polynomial. Finally, we make use of the shape of the sums used to make these estimates and connections with intersection numbers to relate these properties of the gamma vector to algebraic structures (e.g. characteristic classes involved in existing log concavity and Schur positivity properties).
@@ Introduction
This is how it was defined by Gal [21] in the context of to real-rootedness questions of $h$-polynomials of flag spheres and had lower bound results with $h$-vectors in mind related to postivity of the gamma vector. In earlier work of Foata–Schützenberger and Strehl in connection to Eulerian polynomials, the gamma vector appears as the coefficients $\gamma_{i}$ used to write $h(t)$ as a linear combination of $t^{i}(1+t)^{n-2i}$ when $h(t)$ is a reciprocal polynomial. Properties of this vector have since been involved in many different examples in combinatorics with the perspective taken by Gal tying together many objects studied earlier [6].
Since the gamma vector of a polynomial is defined in terms of a functional equation involving the starting polynomial, it is natural to ask if there are explicit expressions in terms of coefficients of the original polynomial. We find explicit formulas for the gamma vector in terms of the coefficients of $h(t)$ in a way that complements the connections to lower bounds and derivatives involving the input polynomial mentioned above. For example, the fact that we get negativity for large ratios (e.g. from considering face vectors as in Corollary 2.2) indicates that the symmetry of the coefficients or the underlying structure of a sphere play a role. In addition, they can be used to extend the definition of the gamma vector to arbitrary polynomials. Concretely, the components of the gamma vector/coefficients of $\gamma(u)$ written as a linear combination of the coefficients of $h(t)$ have coefficients coming from products of Catalan numbers and binomial coefficients (Theorem 1.1). Note that the $n^{\text{th}}$ Catalan number enumerates the number of noncrossing partitions of $\{1,\ldots,n\}$ and that noncrossing partitions and Coxeter groups feature in many examples where gamma positivity is observed (e.g. Coxeter complexes and Eulerian polynomials, Coxeter–Narayana polynomials and the poset of noncrossing partitions [6]). Further details on relations to existing gamma positivity examples and other fields are discussed in Remark 1.8.
@@
Continuing in this direction, we also consider interactions with algebraic structures. Since many of the examples we had considered earlier were connected to independence complexes and broken circuits of matroids (which can be express in terms of intersection numbers), we were motivated to consider an input polynomial with coefficients coming from intersections of divsiors on a variety. So, we considered the case where the input polynomial is the specialization $V(u)=(Au+B)^{d}$ of the volume polynomial associated to a pair of divsiors $A,B$ on a smooth projective variety. In general, the gamma vector has the same sign as $(-1)^{r}$ when the intersections $A^{k}B^{d-k}$ alternate in sign. Suppose that these intersections are nonnegative. When we are working with $\mathbb{P}^{n}$, we can relate the sign of the gamma vector to the relative “size” of $A$ compared to $B$ (since these correspond to integers) (Example 3.4. When the ratio is very large, the gamma vector is negative. If it is very small, it is alternating in sign like the negative ratio case. Positivity properties of divisors are also related to the bounds in a more general setting since the $A^{k}B^{d-k}$ form a log concave sequence when $A$ and $B$ are nef. This can be used to reduce upper and lower bound conditions related to alternating sums of positive multiples of such terms to those on the first or last pair of elements in a sequence. Note that the connections to upper/lower bounds studied in the previous section carry over to volume polynomials with some additional interpretations and simplifications. Finally, we give an explanation of the $\mathbb{P}^{n}$ case in terms of Segre classes of vector bundles and make use of these classes to connect the rational functions giving components of the gamma vector in the derivative expression to positivity questions (e.g. log concavity) related to Segre classes and indicate a general connection between signs of the gamma vector to Schur positivity for nef divisors (Example 3.7).
@@ Acknowledgments
I’d like to thank Lutz Warnke for pointing out a more recent reference [35] describing an elementary approach to the Lagrange inversion formula. Also, I’d like to thank Jon McCammond for mentioning connections between free probability and noncrossing partitions. This work was supported by Horizon Europe ERC Grant number: 101045750 / Project acronym: HodgeGeoComb.
@@ 1.1 An explicit formula involving Catalan numbers, binomial coefficients, and derivatives
where $\widetilde{C}_{k}$ is the $k^{\text{th}}$ Catalan number if $k\geq 1$ and $\widetilde{C}_{0}=0$. Since $\deg\gamma\leq\lfloor\frac{n}{2}\rfloor$ if $h$ is reciprocal, this sum is equal to 00 for $m>\lfloor\frac{n}{2}\rfloor$ in that case.
@@
Remark 1.8.
(Connections to noncrossing partitions)
The Catalan numbers hint at some connections to existing gamma positivity examples. Recall that the $n^{\text{th}}$ Catalan number enumerates the number of noncrossing partitions of $[n]=\{1,\ldots,n\}$. Noncrossing partitions and Coxeter groups feature in many examples connected to gamma positivity (e.g. Eulerian polynomials and $h$-vectors of type A Coxeter complexes, Coxeter–Narayana polynomials and the poset of noncrossing partitions in [34] and Section 2.2 on p. 24 – 27 of [6]). Also, it would be interesting if there is a relation to higher moments in the context of connections between free probability and noncrossing partitions [27], which includes applications ([7], [28], [8]) of the Lagrange inversion formula used to obtain the derivative formula for $r\gamma_{r}$ in Theorem 1.7. Finally, noncrossing partition statistics also feature in an expression for a version of the toric $h$-vector for cubical complexes [9], which are part of the geometric motivation for the study of flag complexes (see discussion on locally CAT(0) spaces on p. 270 of [21]).
@@ 2.1 Bounds involving translations
in order to have $\gamma_{r}\leq 0$ for even $r$ as well. This is a sort of lower bound condition.
@@ 2.2 Upper/lower bounds and general gamma vectors
where we used the identity $\binom{p}{q}-\binom{p-1}{q-1}=\binom{p-1}{q}$ in the last line.
This means that
@@
This follows directly from clearing denominators since $k(r-k-1)=kr-k^{2}-k\leq(k+1)(r-k)=kr-k^{2}+r-k$. The proof of Part 2a) implies that we have shown the statement of Part 2b) for $0\leq r\leq\frac{d}{3}$. To look at the remaining indices, we can split the first sum in Proposition 2.4 into those where $k\geq\frac{r}{d-2r}$ (decreasing coefficients of $b_{k}$) and $k<\frac{r}{d-2r}$ (increasing coefficients of $b_{k}$). The terms that we take alternating sums of are increasing on each part. Lemma 2.5 then shows that the terms have sign $(-1)^{r-1}$ and we have the claimed inequality.